OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -6 and 162, respectively.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

@Jaynator495

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@ikram002p

Nnesha (nnesha):

explicit formula \[\huge\rm a_n = a_1 * r^{n-1}\] where a_1 is first term r =common ratio n=term

Nnesha (nnesha):

for this question we have to find common ratio and first term with given 2nd and 5th \[\large\rm a_2 = -6~,~~~~~~~~~~~~a_5 = 162\] a_2 means 2nd term so when a_2 =-6 n =2 when a_5 = 162 n =5 n here represent the number of terms so we should write two equations \[\huge\rm a_\color{Red}{2 }= a_1 * r^{\color{red}{n}-1}\] a_2 = -6 so i'll replace a_2 with -6 n =2 bec it's a 2nd term \[\huge\rm \color{Red}{-6 }= a_1 * r^{\color{red}{2}-1}\] now write 2nd equation when n=5 a_5 =162

Nnesha (nnesha):

do you understand ?:=))

OpenStudy (anonymous):

could the answer be between these two?

Nnesha (nnesha):

i don't yet do you understand what i said above ?

Nnesha (nnesha):

know*

Nnesha (nnesha):

\[\huge\rm a_n = a_1 * r^{n-1}\] write 2nd equation when n= 5 and a_5 = 162 just like i did for a_2

OpenStudy (anonymous):

162 = a_1 * r^5-1

Nnesha (nnesha):

right 5-1 =4 so it would be \[\huge\rm \color{Red}{162 }= a_1 * r^{\color{red}{5}-1}\] \[\huge\rm \color{Red}{-6 }= a_1 * r^{\color{red}{2}-1}\] first equation \[\large\rm \color{Red}{162 }= a_1 * r^{\color{red}{4}}\] 2nd equation \[\large\rm \color{Red}{-6 }= a_1 * r^{\color{red}{1}}\] i remembered when we solved arithmetic where we subtracted both equation but here we should divide them